Two cyclist are traveling the same distance at the speeds of 10 km/hr and 15 km/hr If one cyclist takes 40 minutes more to cover the distance, the what distance (in Km) are they traveling ?
Answer by mananth(16949)

let distance be x
first cyclist rate = x/10
second cyclist rate = x/15
difference in time = 40 minutes=> 40/60=>2/3 hour
x/10-x/15=2/3
LCD = 30
multiply equation by 30
3x-2x=20
x=20 km is the distance